OpenStudy (anonymous):

A solution of ammonia and water contains 1.80×1025 water molecules and 7.00×1024 ammonia molecules. How many total hydrogen atoms are in this solution

OpenStudy (anonymous):

I'm assuming that it's 10^25 and 10^24 right? @mario95

OpenStudy (anonymous):

yeah im sorry for the typos

OpenStudy (anonymous):

The chemical formula for Ammonia is \(\bf NH_3\) and for water it's \(\bf H_{2}O\). If we have \(\bf 1.8 \times 10^{25}\) water molecules and \(\bf 7.00 \times 10^{24}\) ammonia molecules. To count the number of atoms, we must first dissociate/break up the molecules. Let's start by breaking up ammonia. For each ammonia molecule, we have 3 Hydrogen atoms and 1 Nitrogen atom. Since there is \(\bf 7.00 \times 10^{24}\) ammonia molecules and 3 hydrogen atoms in each molecule, we multiply the number of ammonia molecules by 3 to get the number of hydrogen atoms in the ammonia: \(\bf 3 \times 7.00 \times 10^{24}=2.1 \times 10^{25}\) Hydrogen atoms in ammonia. In each water molecule we have 2 Hydrogen atoms and 1 Oxygen atom. Since there is \(\bf 1.8 \times 10^{25}\) water molecules and 2 Hydrogen atoms in each molecule, we multiply the number of molecules by 2: \(\bf 2 \times 1.8 \times 10^{25}=3.6 \times 10^{25}\) Hydrogen atoms in water. Hence the total number of water atoms in the solution of ammonia and water is:\[\bf \therefore \ Total \ Hydrogen \ atoms = (2.1 \times 10^{25})+(3.6 \times 10^{25})=5.7 \times 10^{25}\]

OpenStudy (anonymous):

@mario95